OpenStudy (anonymous):

Determine whether Mean Value Theorem can be applied to the function on the given interval; if so, find the value(s) of c guaranteed by the theorem. (Enter your answers as a comma-separated list. ) f(x) = x + 1/x on [1, 3]

OpenStudy (anonymous):

yes it is continuous and yes it is differentiable on that interval , so go ahead and find \(\frac{f(3)-f(1)}{3-1}\)

OpenStudy (anonymous):

4/6

OpenStudy (anonymous):

\(f(1)=2,f(3)=\frac{10}{3}, f(3)-f(1)=\frac{10}{3}-1=\frac{7}{3}\)

OpenStudy (anonymous):

oops yes you are right and i am wrong

OpenStudy (anonymous):

\(\frac{10}{3}-2=\frac{4}{3}\) divide by 2 get \(\frac{2}{3}\) take the derivative, set it equal to \(\frac{2}{3}\) and solve for \(x\)

OpenStudy (anonymous):

i.e. \(1-\frac{1}{x^2}=\frac{2}{3}\) etc

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but there's more to the problem